OpenStudy (anonymous):

Two trinomials are multiplied together. If the degree of the first trinomial is 4, what must the degree of the second trinomial be in order to have a product whose degree is 8?

OpenStudy (campbell_st):

the degree of the 2nd can be found by using the index law for multiplication if you multiply the leading terms \[x^4 \times x^? = x^8\] remember... the index law says add the powers